OpenStudy (anonymous):

Factor 4sˆ5+16s³-32s-20 4s^5 is suppose to be 4s exponent5 but i dont know how to do it

OpenStudy (anonymous):

What have you done so far?

OpenStudy (anonymous):

nothing

OpenStudy (anonymous):

Well, do you know how to find the factors of each term?

OpenStudy (anonymous):

cf is 4

OpenStudy (anonymous):

That is a start. Pull a factor of 4 from each term and see what you have left.

OpenStudy (anonymous):

4, 8, 5

OpenStudy (anonymous):

No, I mean write the new expression after you factor out a 4

OpenStudy (anonymous):

4(4s^3-8s-5)

OpenStudy (anonymous):

Not quite. You're missing a s^5, but otherwise that's good.

OpenStudy (anonymous):

Now what other factors do some of your terms have in common?

OpenStudy (anonymous):

s

OpenStudy (anonymous):

Ok, so factor those out from the terms that have it

OpenStudy (anonymous):

not sure how

OpenStudy (anonymous):

show me step by step plz and then explain

OpenStudy (anonymous):

How did you factor out the 4?

OpenStudy (anonymous):

4(4s^3-8s-5)

OpenStudy (anonymous):

Right, but I mean what did you do?

OpenStudy (anonymous):

Also that's not right

OpenStudy (anonymous):

divide each term by 4

OpenStudy (anonymous):

You keep forgetting the 4s^5

OpenStudy (anonymous):

then what would it be

OpenStudy (anonymous):

4(s^5 +4s^3 - 8s - 5)

OpenStudy (anonymous):

But as I said, you can factor an s from the first 3 terms inside the parens.

OpenStudy (anonymous):

Or rather, like you said.

OpenStudy (anonymous):

And you would do it the same way. Put an s out front, and divide each term by the s you took away.

OpenStudy (anonymous):

Just don't take one away from the term that doesn't have an s to start with.

OpenStudy (anonymous):

s(^5+4^3-8-5)

OpenStudy (anonymous):

No. What is \[\frac{s^5}{s}\]

OpenStudy (anonymous):

idk im really bad with fractions

OpenStudy (anonymous):

Then you should practice! =) Everything depends on everything else. If you practice your fractions until you are not bad with them, factoring will be a very fast process.

OpenStudy (anonymous):

Lets start now. What is \[\frac{32}{2}\]

OpenStudy (anonymous):

16

OpenStudy (anonymous):

Ok, and if I said \[\frac{2^5}{2}\] ?

OpenStudy (anonymous):

rawr!

OpenStudy (anonymous):

a^5

OpenStudy (anonymous):

What is \(2^5\)

OpenStudy (anonymous):

32

OpenStudy (anonymous):

Interesting. And what was \[\frac{32}{2} \] again?

OpenStudy (anonymous):

16/1

OpenStudy (anonymous):

Right.. Just curious, but what is 2^4?

OpenStudy (anonymous):

rawr! rawr!

OpenStudy (anonymous):

16

OpenStudy (anonymous):

That's interesting. \[\frac{2^5}{2} = 2^4\]

OpenStudy (anonymous):

Does this work with other numbers? \[3^3 = 27\] \[\frac{27}{3} =\ ?\]

OpenStudy (anonymous):

9

OpenStudy (anonymous):

And 9 is \(3^2\)

OpenStudy (anonymous):

Lets see if we can figure out why this might be happening.

OpenStudy (anonymous):

What is \(a^2\)

OpenStudy (anonymous):

Or better still, \(a^3\)

OpenStudy (anonymous):

a*a*a

OpenStudy (anonymous):

Ok, now if we have \[\frac{a\cdot a\cdot a}{a}\] We can cancel one of the a's on top with the a on the bottom. So what would that equal?

OpenStudy (anonymous):

a^2

OpenStudy (anonymous):

Right. So now, coming back to the original question.. What is \[\frac{s^5}{s}\]

OpenStudy (anonymous):

5

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

then just s

OpenStudy (anonymous):

What was \[\frac{a^3}{a}\] again?

OpenStudy (anonymous):

a*a*a

OpenStudy (anonymous):

No, you just said it 6 mins ago. just scroll up

OpenStudy (anonymous):

a^2

OpenStudy (anonymous):

And do you remember why?

OpenStudy (anonymous):

because we cancel out 1

OpenStudy (anonymous):

So now. What is \[\frac{s^5}{s}\] Use the same argument as before

OpenStudy (anonymous):

s^4

OpenStudy (anonymous):

yes, precisely.

OpenStudy (anonymous):

So if we divide s^5 by s, we get s^4

OpenStudy (anonymous):

And if we divide s^3 by s?

OpenStudy (anonymous):

s^2

OpenStudy (anonymous):

And if we divide s by s?

OpenStudy (anonymous):

s

OpenStudy (anonymous):

s/s = s?

OpenStudy (anonymous):

What is 2/2

OpenStudy (anonymous):

1

OpenStudy (anonymous):

What is 5/5

OpenStudy (anonymous):

1

OpenStudy (anonymous):

And w/w?

OpenStudy (anonymous):

w

OpenStudy (anonymous):

Why was 5/5 = 1?

OpenStudy (anonymous):

s/s=1

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

s goes into s 1 time

OpenStudy (anonymous):

that's a good reason.

OpenStudy (anonymous):

So now.

OpenStudy (anonymous):

thats why i said s i thought instead of putting 1 you put the variable

OpenStudy (anonymous):

If we need to factor an s from s^5 +4s^3 - 8s what will our expression be?

OpenStudy (anonymous):

s^4+s^2-8

OpenStudy (anonymous):

not quite.

OpenStudy (anonymous):

But very close!

OpenStudy (anonymous):

i hate this math

OpenStudy (anonymous):

\[s^5 + 4s^3 - 8s\]

OpenStudy (anonymous):

factor out an s, does that effect the numbers out in front?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

so where'd the 4 go when you factored?

OpenStudy (anonymous):

s(s^4+4s^2-8-5)?

OpenStudy (anonymous):

Where'd that 5 come from?

OpenStudy (anonymous):

the original equation

OpenStudy (anonymous):

jumped the gun a bit

OpenStudy (anonymous):

That 5 didn't have an s